# Funnel Synthesis via LMI Copositivity Conditions for Nonlinear Systems

### v2

@@ 5 Conclusion
This paper presents a funnel synthesis algorithm for nonlinear systems under bounded disturbances. The proposed method can satisfy the invariance condition exactly over the finite time horizon. This is achieved by approximating the incremental system as an uncertain LPV system and deriving a finite number of LMIs that imply the DLMI resulting from the invariance condition. Simulation results show that the synthesized funnel satisfies both invariant and feasibility conditions.
@@ Acknowledgement
This work was supported in part by the Air Force Office of Scientific Research under Grant FA9550-20-1-0053 and in part by the Office of Naval Research under Grant ONR N00014-20-1-2288.

### v3

@@ 5 Conclusion
This paper presents a funnel synthesis algorithm for nonlinear systems under bounded disturbances. The proposed method can satisfy the invariance condition exactly over the finite time horizon. This is achieved by approximating the incremental system as an uncertain LPV system and deriving a finite number of LMIs that imply the DLMI resulting from the invariance condition. Simulation results show that the synthesized funnel satisfies both invariant and feasibility conditions. Future work will focus on extending the proposed method to continuous-time constraint satisfaction, not only for the invariance condition but also for state and input constraints. Improving computational efficiency for large-scale SDP problems is also a valuable direction.
@@ Acknowledgement
This work was supported in part by the Air Force Office of Scientific Research under Grant FA9550-20-1-0053 and in part by the Office of Naval Research under Grant ONR N000142512231.